Two helical springs of the same material ( permissile shear stress is 5800 KN/cm^2 , and modulus of rigidity is 10000 KN/m^2 ) and free length are used to make a concentric spring. The concentric spring is compressed 2.8 cm by an axial load of 360 KN. Assuming that spring index of both springs is 7 foe each spring .


Determine for each spring :

load ,

stored energy ,

angular deflection ,

number of coils ,

wire diameter ,

spring diameter .



1
Expert's answer
2021-03-12T05:42:29-0500

"C = \dfrac{D}{d}"


7 = D/2.8

D = 19.6 cm

spring diameter = 19.6cm



"\theta = \dfrac{16FD²}{Gd^4} ="


"\dfrac{16×360×19.6² }{5800×2.8^4} ="


6.2 rad